OpenStudy (anonymous):

The volume of a rectangular prism is 115 cm3. If the length is doubled, the width is tripled, and the height is cut in half, what is the volume of the new prism?

OpenStudy (cwrw238):

volume = l*w*h = 115 new volume = 2l*3w*0.5h dividing 2l*3w*0.5h / l*w*h = 2*3*0.5 = 3 volume of new prism = 115 * 3 = 345

OpenStudy (cwrw238):

cm3

OpenStudy (anonymous):

thankyou

OpenStudy (cwrw238):

yw